OpenStudy (anonymous):

Find the solutions 18cot2x - 6 = 0 in the interval from 0 to inclusive. ??

OpenStudy (anonymous):

so 18. cot 2x -6 = 0 Or 18 cot 2x = 6 Now tell me if you meant cot 2x or \[\cot ^2x\]

OpenStudy (anonymous):

Whoops, cot^2 x

OpenStudy (anonymous):

Then its simple. you know that cot^2x now is 6/18 i.e. cot^2x = 1/3. You can take it from here?

OpenStudy (anonymous):

@cmolina19 tell me this; have you studied about trigonometric ratios of standard angles like 30,60, etc?

OpenStudy (anonymous):

well it is on this lesson. I have a table with the common first quadrant angles, but that's about it. Kind of lost though :/

OpenStudy (anonymous):

Ok don't worry. http://www.schooltutoring.com/help/trigonometric-ratios-of-standard-angles/ This is the table of trigonometric values of angles of 0,30,45,60,90. Coming to your question, cot^2x = 1/3 Hence, cot x = 1/sq. root of 3 (taking sq. root of both sides) (Refer to the table, and find out for what angle the value of cot x = 1/ sq. root of 3) Lets say cot 20 is 1/sq. root of 3 (although its not) Then we can say, since cot x = 1 / sq. root of 3 cot x = cot 20 Hence x = 20. (Actually the answer is NOT 20; i want you to find the real value of x)

OpenStudy (anonymous):

Gotcha. Thanks. :b

OpenStudy (anonymous):

anyway the ans is 60. :P

OpenStudy (anonymous):

yeah xD Took a while since I had to stare at the table for a while, needed my answer in radians. Shoulda specified.

OpenStudy (anonymous):

Oh if its radians then the answer is Pi over 3 . cheers! :)